The length of the base of a triangle is twice its height. If the area of the triangle is 196 square​ kilometers, find the height.
Answer by rothauserc(4718)

Area(A) of a triangle = (1/2) * base * height
:
let x be the length of the height, then base = 2x
:
we are given A = 196 square kilometers
:
196 = (1/2) * 2x * x
:
x^2 = 196
:
x = square root(196) = + or - 14 kilometers
:
********************************************************
height of triangle = 14 kilometers
:
Note that we reject the negative value(-14) for the base
********************************************************
:
we check the answer
:
196 = (1/2) * 28 * 14 = 196
:
our answer checks
: